Recording events in a horizontal statements model The following events pertain to Super Cleaning Company:

1. Acquired $10,000 cash from the issue of common stock. 2. Provided $15,000 of services on account. 3. Provided services for $5,000 cash. 4. Received $2,800 cash in advance for services to be performed in the future. 5. Collected $12,200 cash from the account receivable created in Event 2. 6. Paid $1,900 for cash expenses. 7. Performed $1,400 of the services agreed to in Event 4. 8. Incurred $3,600 of expenses on account. 9. Paid $4,800 cash in advance for one-year contract to rent office space. 10. Paid $2,800 cash on the account payable created in Event 8. 11. Paid a $1,500 cash dividend to the stockholders. 12. Recognized rent expense for nine months’ use of office space acquired in Event 9 Required Show the effects of the events on the financial statements using a horizontal statements model like the following one. In the Cash Flows column, use the letters OA to designate operating activity, IA for investing activity, FA for financing activity, and NC for net change in cash. Use NA to indicate accounts not affected by the event. The first event is recorded as an example.

Effect of Events on the Financial Statements

Balance Sheet

Income Statement

Stmt. of

Assets

=

Liabilities

+

Stockholders’ Equity

Rev.

-

Exp.

=

Net Inc.

Cash Flows

Event

No.

Cash

+

Accts Rec.

+

Pp. Rent

=

Accts. Pay.

+

Unearn Rev.

+

Com. Stock

+

Ret. Earn.

1.

10,000

+

NA

+

NA

=

NA

+

NA

+

10,000

+

NA

NA

-

NA

=

NA

10,000   FA

2.

+

+

=

+

+

+

-

=

3.

+

+

=

+

+

+

-

=

4.

+

+

=

+

+

+

-

=

5.

+

+

=

+

+

+

-

=

6.

+

+

=

+

+

+

-

=

7.

+

+

=

+

+

+

-

=

8.

+

+

=

+

+

+

-

=

9.

+

+

=

+

+

+

-

=

10.

+

+

=

+

+

+

-

=

11.

+

+

=

+

+

+

-

=

12.

+

+

                  *

=

+

+

+

-

=

Bal.

$                 

+

$             

=

$             

+

$             

+

$             

$             

-

$             

$             

$             

=

*$          x    /12 = $
Answer #1
Effect of Events on the Financial Statements
Balance Sheet Income Statement Stmt. of
Assets = Liabilities + Stockholders’ Equity Rev. - Exp. = Net Inc. Cash Flows
Event Cash + Accts Rec. + Pp. Rent = Accts. Pay. + Unearn Rev. + Com. Stock + Ret. Earn.
No.
1 10,000 + NA + NA = NA + NA + 10,000 + NA NA - NA = NA 10,000   FA
2 + 15000 + = + + + 15000 15000 - = 15000
3 5000 + + = + + + 5000 5000 - = 5000 5000
OA
4 2800 + + = + 2800 + + - = 2800
OA
5 12200 + -12200 + = + + + - = 12200
OA
6 -1900 + + = + + + -1900 - 1900 = -1900 -1900
OA
7 + + = + -1400 + + 1400 1400 - = 1400
8 + + = 3600 + + + -3600 - 3600 = -3600
9 -4800 + + 4800 = + + + - = -4800
OA
10 -2800 + + = -2800 + + + - = -2800
OA
11 -1500 + + = + + + -1500 - = -1500
FA
12 + + -3600 = + + + -3600 - 3600 = -3600
Bal. 19,000 + 2,800 = 19000 + 800 + 1,400 10,000 - 10,800 21,400 9,100 = 12,300 19,000
NC
$4,800 X 9/12 = $3,600